OpenStudy (anonymous):

is this correct? using l'hopitals rule limit as x goes to 1+ ln(x^2)/(x^2-1)=1

OpenStudy (anonymous):

first off start with \[\frac{2\ln(x)}{x^2-1}\]

OpenStudy (anonymous):

this will make your differentiation much easier then you should be good to go

OpenStudy (anonymous):

First confirm that it's indeterminate form. Evaluating at x=1 gives 0/0. Cool, we can use L'hopital's! Take the derivative on top and on bottom gives: \[\Large \frac{\frac{1}{x^2}*2x}{2x}\]

OpenStudy (anonymous):

Any questions on that?

OpenStudy (anonymous):

one i understand what you did

OpenStudy (anonymous):

what smoothmath said then to a tiny big of algebra, you will see that the limit is not 1

OpenStudy (anonymous):

none*

OpenStudy (anonymous):

Isn't it though, Satellite?

OpenStudy (anonymous):

Cool, if you understand those steps, then all you have to do is evaluate at x=1, which gives you 2/2 =1

OpenStudy (anonymous):

So yeah, you had the right answer.

OpenStudy (anonymous):

yes all good :) thanks

OpenStudy (anonymous):

Unless I made some mistake or Satellite is seeing something I don't.

OpenStudy (anonymous):

I dont know but that would be 2 to 1

OpenStudy (anonymous):

My pleasure! =D

OpenStudy (anonymous):

oh yes, it is one for an idiotic moment i was replacing x by 0!